Robert L. Krawitz

428 Cambridge St. Cambridge, MA 02141

(617) 661-2538

Objective A position in software engineering with an aim toward project management.

Education  Massachusetts Institute of Technology  1982-1987

S.B., Computer Science (June 1987) Courses taken include Software Engineering Laboratory, Computer Systems Engineering, Introduction to Algorithms, Contemporary Computer Engineering; Humanities Concentration in Economics. Thesis: development of flexible data communications system for MIT Center for Meteorology and Physical Oceanography.

Experience  Thinking Machines Corporation  Cambridge, MA

Member of Technical Staff June 1987—present Member of Numerical and Scientific Computation Group (NCG), and Research Computing Group. In the NCG: engineering/maintenance of elementary functions library for Connection Machine (CM-2); reengineering of Fast Fourier transform. Currently working on the Connection Machine Scientific Software Library, a library of basic scientific subroutines based upon a common data abstraction. Performance analysis work using sample applications. In the Research Computing Group: maintenance of internal network of several hundred Symbolics Lisp Machines, VAXes, Sun workstations, and other assorted hardware. Designing and implementing plan to streamline the configuration of the network to improve availability and maintainability of resources based upon a “standard server” model.

MIT Project Athena Cambridge, MA Systems Programmer September 1985—May 1987 Development of system software for approximately 50 DEC VAX minicomputers, and many VAXstation II workstations and IBM RT-PC workstations running 4.3 BSD UNIX. Projects include maintenance and extension of GNU Emacs text editor and development of utilities. Operations Staff June 1985—August 1985 Routine system management of DEC VAX 11/750 computers. Increasing involvement with Systems Development Group throughout period.

ION Weather/Jersey Weather Center 1978—1984 Weather observations, interpretation of weather data, and related tasks for commercial weather service. Developed microcomputer system for continuous collection and processing of international weather data for clients.

Activities MIT Men’s Basketball Team Manager 1982—1986 Head manager 1984—1986. Responsible for coordinating activities of managerial staff of three, and members of team. Coordinator of Operations, David Koch Classic, a basketball tournament involving eight teams, November 1986. Author of Managers’ Manual, a guide for intercollegiate team managers at MIT, 1985 (revised 1986). Burton R. Anderson award for Manager of the Year, 1986.

Other activities included Student Information Processing Board, a student group that encourages use of computers by students on campus; Class of 1986 Council.

Moderator of Internet interest group “info-nets” since 1984.

Publications GNU Emacs Programming (lecture notes). Available from SIPB at MIT. Contributions to GNU Emacs Programmer’s Manual (currently unpublished).

Background Operating Systems: Unix BSD 4.2, 4.3; SunOS 3.x, 4.0; Ultrix; Symbolics Genera 7.1, 7.2

Languages: C, Common Lisp, GNU Emacs Lisp, CLU, Connection Machine Languages (*Lisp, C*, Paris), Z-80 assembly language.

Hardware: Sun 3, Sun 4, VAX, Connection Machine, RT/PC.